### v3

@@ 3.1 Existing Approaches
A common strategy used to estimate $\theta_{I}^{*}$, for example, is to compute the difference in means between control units in treated and untreated clusters. Letting $c(i)\in\{1,\ldots,k\}$ denote the index of the cluster containing unit $i$, this corresponds to
$\hat{\theta}_{I}^{+}=\frac{\sum_{i\in\mathcal{N}_{n}}(1-D_{i})W_{c(i)}Y_{i}}{\sum_{i\in\mathcal{N}_{n}}(1-D_{i})W_{c(i)}}-\frac{\sum_{i\in\mathcal{N}_{n}}(1-W_{c(i)})Y_{i}}{\sum_{i\in\mathcal{N}_{n}}(1-W_{c(i)})},$ (1)
which is the sample analog of
$\bar{\theta}_{I}^{+}=\frac{1}{n}\sum_{i\in\mathcal{N}_{n}}\big{(}{\bf E}[Y_{i}\mid(1-D_{i})W_{c(i)}=1]-{\bf E}[Y_{i}\mid W_{c(i)}=0]\big{)}.$
The problem is that $\bar{\theta}_{I}^{+}$ may be quite far from the actual target $\theta_{I}^{*}$. For instance, units in control clusters near cluster boundaries may be spatially proximate to units in treated clusters and therefore at greater risk of contamination. For such units $i$, ${\bf E}[Y_{i}\mid W_{c(i)}=0]$ may be substantially different from $Y_{i}(\bm{0})$.
The fried-egg design attempts to reduce bias by restricting the sums in the formula to the subset of units sufficiently far from cluster boundaries. Unfortunately, this has two problems. First, the resulting estimator is actually asymptotically biased because boundary units are excluded with probability one, so it only estimates an average effect for the subpopulation of units in cluster interiors. As noted by McCann et al. (2018), this differs from the target estimand since units in cluster interiors may be systematically different from those near the boundaries due to spatial heterogeneity. Second, it is inefficient. The purpose of only including units in the interiors of control clusters is that such units are “well surrounded” by control clusters, or equivalently, relatively far from treated clusters. However, boundary units may be well surrounded in the same fashion if all neighboring clusters are assigned to control, so it would be just as useful to include these units.
@@ 3.2 Our Approach
that is, if a unit $i$’s $r_{n}$-neighborhood only intersects clusters assigned to the same treatment arm. Our estimation strategy is to exclude from estimation units not well surrounded. If $r_{n}=0$, then all units are well surrounded, and our estimators will reduce to difference in means. Choosing a larger radius $r_{n}$ is analogous to choosing a larger boundary region for exclusion in a fried-egg design.
Our strategy excludes strictly fewer units than the fried-egg design since some boundary units will be well surrounded. It also solves the fried-egg design’s bias problem because any unit has nontrivial probability of being well surrounded. In other words, it avoids positivity violations or limited overlap (see Remark 1 below), whereas boundary units have zero overlap under the fried-egg design. Boundary units under our scheme are less likely to be well surrounded than interior units. However, we account for this using inverse-probability weighting, and we show formally that the number of units that are well surrounded is of asymptotic order equal to the population size $n$. Figure 1 depicts $k$-medoid clusters, marking with an “X” units excluded under our scheme due to not being well surrounded.
@@
For $d_{t}\in\{0,1\}$, define $T_{ti}=\bm{1}\{D_{i}=d_{t},W_{c(i)}=t\}S_{i}$, and for $d_{t}=\varnothing$, $T_{ti}=\bm{1}\{W_{c(i)}=t\}S_{i}$. Let $p_{ti}={\bf E}[T_{ti}]$. We propose the following Hájek estimator for $\theta^{*}$:
@@
Remark 1 provides a closed-form expression for the propensity score $p_{ti}$. When $r_{n}=0$, $S_{i}=1$ for all $i$, so propensity scores are homogeneous across $i$, and $\hat{\theta}$ reduces to difference in means. For $r_{n}>0$, the scores are generally heterogeneous.
We next specify the radius $r_{n}$. For any cluster $C_{j}$, let $m_{j}$ be its centroid from Assumption 5 and $R_{j}=\max_{i\in C_{j}}\rho(i,m_{j})$, the “radius” of the cluster. We set
@@
where $p_{t}^{\varnothing}\equiv 1$. We show in Lemma SA.5.3 that $\phi_{i}$ is asymptotically bounded uniformly in $i$ under (2). This implies that the number of units that is well surrounded is of asymptotic order equal to the population size $n$. Given that $q\in(0,1)$ by Assumption 4, $p_{ti}$ is similarly bounded away from zero under appropriate conditions on the saturation levels $p_{1},p_{0}$ (see Assumption 6 below).
@@
The choice of 0.5 in (2) is immaterial for the asymptotic theory, but in finite samples it controls a bias-variance trade-off. We choose 0.5 to informally balance the two. Constants close to zero will yield high bias because $r_{n}=0$ corresponds to difference in means. Constants close to one result in high variance since most units are excluded with high probability. To see this, suppose each cluster is a ball with homogeneous radius $R$. Setting the constant to 1 implies $r_{n}=R$, so the only unit in any cluster that is well surrounded with probability one is the centroid. A unit located slightly north of the centroid finds its $r_{n}$-neighborhood intersecting at least two clusters, so the chance that it is well surrounded is substantially lower. In contrast, suppose $r_{n}=0.5R$, and define a cluster $C_{j}$’s “interior” as the ball centered at $m_{j}$ with radius $R/2$. Then all units in the interior are well surrounded with probability one because their $r_{n}$-neighborhoods are contained within $C_{j}$.
@@
$\Lambda_{i}=\big{\{}\ell\in\mathcal{N}_{n}\colon\exists\,j\,\text{ s.t. }\,C_{j}\cap\mathcal{N}(i,r_{n})\neq\varnothing,\,C_{j}\cap\mathcal{N}(\ell,r_{n})\neq\varnothing\big{\}},$ (3)
the set of units whose $r_{n}$-neighborhoods intersect a cluster that intersects $i$’s $r_{n}$-neighborhood. These can be thought of as the units “most potentially correlated” with $i$. Define $A_{ij}(1)=\bm{1}\{j\in\Lambda_{i}\}$, $A_{ij}(2)=\bm{1}\{j\in C_{c(i)}\}$, and $\hat{Z}_{i}=(T_{1i}(Y_{i}-\hat{\mu}_{1}))/p_{1i}-(T_{0i}(Y_{i}-\hat{\mu}_{0}))/p_{0i}$. The variance estimator is
@@
Notice that $\hat{\sigma}^{2}(2)$ is the conventional cluster-robust variance estimator (e.g. Baird et al., 2018), which only accounts for within-cluster dependence. The estimator $\hat{\sigma}^{2}(1)$ is analogous to that of Leung (2022b), and it additionally accounts for cross-cluster dependence since $C_{c(i)}\subseteq\Lambda_{i}$. In § 4, we discuss the advantage of taking the larger of the two estimators. We can construct a 95-percent confidence intervals (CIs) as follows:
$\hat{\theta}\pm 1.96\cdot\hat{\sigma}k^{-1/2}.$ (5)
@@ 4 Asymptotic Theory
We derive bounds on the rates of convergence of our estimator and difference in means. Then we characterize the asymptotic distribution of our estimator and prove that the variance estimator is asymptotically conservative. Recall the asymptotic order notation from the end of § 1. Further recall from § 2.3 that the estimand is $\theta^{*}\equiv\theta^{*}(d_{1},d_{0},p_{1},p_{0})$ for $d_{1},d_{0}\in\{\varnothing,0,1\}$ and $p_{1},p_{0}\in[0,1]$.
@@
For any $t\in\{0,1\}$, if $d_{t}=1$ ($d_{t}=0$), then $p_{t}>0$ ($p_{t}<1$).
As discussed in Remark 1, this condition is sufficient to ensure overlap or positivity.
@@
The result establishes a bias-variance trade-off in $k_{n}$. The $k_{n}^{-1/2}$ term is the contribution of (the square root of) the variance since the effective sample size in a CRT is the number of clusters $k_{n}$. The asymptotic bias is order $r_{n}^{-\gamma}$. Intuitively, if $r_{n}$ is large, then the expected outcome of a unit $i$ assigned treatment $d_{t}$ and well surrounded by clusters assigned to arm $t$ should well approximate ${\bf E}^{*}_{p_{t}}[Y_{i}(d_{t},\bm{D}_{-i})]$, corresponding to lower bias. The bias decreases at a faster rate for larger $\gamma$ since this corresponds to a lower degree of spatial interference.
The requirement $k_{n}\precsim n/\xi_{n}$ is mild, and typically we would have $k_{n}\prec n/\xi_{n}$. In the increasing-domain case where $\xi_{n}=1$, necessarily $k_{n}\precsim n/\xi_{n}$ since $k_{n}\leq n$, but usually the number of clusters is smaller order than the population size. In the infill-increasing case, if $k_{n}$ grows faster than $n/\xi_{n}$, which is the volume of the study region under Assumption 2, then we would be increasingly subdividing the region into smaller clusters of shrinking volume, analogous to having more clusters than units.
Denote the difference-in-means estimator by $\hat{\theta}^{+}$, which corresponds to setting $r_{n}=0$ in the definition of $\hat{\theta}$. Let $\hat{\theta}_{D}^{+}$, $\hat{\theta}_{T}^{+}$, and $\hat{\theta}_{O}^{+}$ be the difference-in-means estimates of $\theta_{D}^{*}$, $\theta_{T}^{*}$, and $\theta_{O}^{*}$.
@@
(a) Under Assumptions 16, $\lvert\hat{\theta}^{+}-\theta^{*}\rvert\precsim(k_{n}\xi_{n}/n)^{1/d}+k_{n}^{-1/2}$. (b) Suppose the design satisfies Assumptions 4 and 5 for $p_{1}\in(0,1)$, $p_{0}=0$. There exists a metric space and sequence of potential outcomes satisfying Assumptions 13 such that $\lvert\hat{\theta}_{Q}^{+}-\theta_{Q}^{*}\rvert\succsim(k_{n}\xi_{n}/n)^{1/d}+k_{n}^{-1/2}$ for any $Q\in\{D,T,O\}$.
Part (a) provides an upper bound on the rate. Part (b) shows that the rate is tight for several illustrative cases. Recall that setting $p_{0}=0$ corresponds to the estimands defined in Hayes and Moulton (2017) where the baseline is “pure control.”
The $k_{n}^{-1/2}$ term in the bound is (the square root of) the variance contribution, as in Theorem 2, while $(k_{n}\xi_{n}/n)^{1/d}$ is the bias contribution, which is notably worse than of $\hat{\theta}_{Q}$. This implies that even a large decrease in the degree of interference $1/\gamma$ can have no effect on the asymptotic order of the bias. The reason is that units situated at cluster boundaries may be directly proximate to clusters assigned to different treatment arms, and as shown in the proof, the share of such units can be of order $(k_{n}\xi_{n}/n)^{1/d}$.
These results demonstrate that excluding units in the manner of $\hat{\theta}$ can significantly reduce the asymptotic order of the bias relative to difference in means. Furthermore, while excluding units may come at the cost of efficiency, there is no increase to the asymptotic order of the variance, so in this sense, the efficiency loss is second order relative to the bias reduction. The reason is because the variances of both estimators scale not with the number of units but with the number of clusters.
@@
The proof of Theorem 2 demonstrates that $\sigma_{n}^{2}\precsim 1$.
@@
The first result (6) centers the estimator at $\bar{\theta}$, the probability limit of $\hat{\theta}$. This is not the estimand of interest since there is an asymptotic bias $\lvert\bar{\theta}-\theta^{*}\rvert$ by Theorem 2. To use the normal limit to justify the validity of the CI (5), the second result (7) considers “undersmoothed designs” in which the number of clusters is of smaller order than the optimal rate discussed in the next section. Here the asymptotic bias is small, so (7) can center at the estimand $\theta^{*}$. This is analogous to nonparametric regression where rate-optimal tuning parameter choices result in asymptotic bias, so use of the normal approximation requires undersmoothing. We discuss choosing $k_{n}$ to undersmooth in § 5. An alternative to undersmoothing is “bias-aware” inference, which we discuss in § SA.2.3.
@@
Suppose $1\prec k_{n}\prec n/\xi_{n}$ and $\sigma_{n}^{2}\succsim 1$. Under Assumptions 16, $\hat{\sigma}^{2}(u)=\sigma_{n}^{2}+\mathcal{B}_{n}+o_{p}(1)$ for $t\in\{1,2\}$ and some sequence of non-negative constants $\{\mathcal{B}_{n}\}_{n\in\mathbb{N}}$. Hence, both $\hat{\sigma}^{2}(1)$ and $\hat{\sigma}^{2}(2)$ are asymptotically conservative.
The cluster-robust variance estimator $\hat{\sigma}^{2}(2)$ has the advantage of being non-negative in finite sample, unlike $\hat{\sigma}^{2}(1)$ which is a truncation estimator (Andrews, 1991, p. 823). On the other hand, $\hat{\sigma}^{2}(2)$ only accounts for within-cluster dependence, whereas $\hat{\sigma}^{2}(1)$ also accounts for cross-cluster dependence in the definition $A_{ij}(1)$. Nonetheless, $\hat{\sigma}^{2}(2)$ is valid because when $k_{n}\rightarrow\infty$, dependence across clusters is second order. However, as shown in the proof, this dependence vanishes very slowly at the rate $(k_{n}\xi_{n}/n)^{1/d}$, the same order as the bias of difference in means. Thus in smaller samples, ignoring second-order terms may result in anti-conservativeness. By taking the larger of the two estimators, we obtain the benefits of both. In § SA.3 we relate this to the literature on variance estimation in settings with interference.

### v4

@@ 3.1 Existing Approaches
Let $c(i)\in\{1,\ldots,k\}$ denote the index of the cluster containing unit $i$. A common strategy for estimating $\theta_{I}^{*}$ is to compute the difference in means between control units in clusters assigned saturation level $p_{1}$ and those in clusters assigned level $p_{0}$:
$\frac{\sum_{i\in\mathcal{N}_{n}}(1-D_{i})W_{c(i)}Y_{i}}{\sum_{i\in\mathcal{N}_{n}}(1-D_{i})W_{c(i)}}-\frac{\sum_{i\in\mathcal{N}_{n}}(1-W_{c(i)})Y_{i}}{\sum_{i\in\mathcal{N}_{n}}(1-W_{c(i)})}.$ (1)
This is the sample analog of
$\frac{1}{n}\sum_{i\in\mathcal{N}_{n}}\big{(}{\bf E}[Y_{i}\mid(1-D_{i})W_{c(i)}=1]-{\bf E}[Y_{i}\mid W_{c(i)}=0]\big{)},$
which may be quite far from the target $\theta_{I}^{*}$. Units in control clusters near cluster boundaries may be spatially proximate to units in treated clusters and therefore at greater risk of contamination. For such units $i$, ${\bf E}[Y_{i}\mid W_{c(i)}=0]$ may be substantially different from ${\bf E}^{*}_{p_{0}}[Y_{i}(0,\bm{D}_{-i})]$.
The fried-egg design attempts to reduce bias by restricting the comparison in (1) to the subset of units deemed sufficiently far from cluster boundaries. Unfortunately, this has two problems. First, the resulting estimator is in fact asymptotically biased because boundary units are excluded with probability one, so it only estimates an average effect for the subpopulation of units in cluster interiors. As noted by McCann et al. (2018), this differs from the target estimand since units in cluster interiors may be systematically different from those near the boundaries due to spatial heterogeneity. Second, it is inefficient. The purpose of only including units in the interiors of control clusters is that such units are “well surrounded” by control clusters, or equivalently, relatively far from treated clusters. However, boundary units may be well surrounded in the same fashion if all neighboring clusters are assigned to control, so it would be just as useful to include these units.
@@ 3.2 Our Approach
that is, if a unit $i$’s $r_{n}$-neighborhood only intersects clusters assigned to the same treatment arm. Figure 1 depicts $k$-medoid clusters, marking with an “X” units that are not well surrounded. Our strategy is to only exclude from estimation units that are not well surrounded. If $r_{n}=0$, then all units are well surrounded, and our estimators reduce to difference in means. Choosing a larger radius $r_{n}$ is analogous to choosing a larger boundary region for exclusion in a fried-egg design.
@@
Under our assumptions, the number of well-surrounded units is of asymptotic order equal to the population size $n$ because any unit has nontrivial probability of being well-surrounded (Lemma SA.5.3). As a consequence, our strategy solves the fried-egg design’s boundary bias problem and typically excludes strictly fewer units.
Let $d_{1},d_{0}$ be given from the estimand $\theta^{*}$. For any $t\in\{0,1\}$ and $i\in\mathcal{N}_{n}$, define $T_{ti}=\bm{1}\{D_{i}=d_{t},W_{c(i)}=t\}S_{i}$ if $d_{t}\in\{0,1\}$ and $T_{ti}=\bm{1}\{W_{c(i)}=t\}S_{i}$ if $d_{t}=\varnothing$. Let $p_{ti}={\bf E}[T_{ti}]$ be the propensity score, which has a closed-form expression given in Remark 1 below. We propose the following Hájek estimator for $\theta^{*}$:
@@
When $r_{n}=0$, $S_{i}=1$ for all $i$, so propensity scores are homogeneous across $i$, and $\hat{\theta}$ reduces to difference in means. For $r_{n}>0$, the scores are generally spatially heterogeneous. For instance, boundary units are less likely to be well surrounded than interior units because that would require more clusters to be assigned the same saturation level.
For any cluster $C_{j}$, let $m_{j}$ be its centroid from Assumption 5 and $R_{j}=\max_{i\in C_{j}}\rho(i,m_{j})$, the “radius” of the cluster. We propose setting
@@
where $p_{t}^{\varnothing}\equiv 1$. We show in Lemma SA.5.3 that $\phi_{i}$ is asymptotically bounded uniformly in $i$ under (2). Since $q\in(0,1)$ by Assumption 4, $p_{ti}$ is similarly bounded away from zero when the saturation levels $p_{1},p_{0}$ are nontrivial (see Assumption 6 below). Hence, using a CRT paired with our estimation strategy inherently avoids positivity violations or limited overlap.
@@
The choice of 0.5 in (2) is immaterial for the asymptotic theory, but in finite samples it controls a bias-variance trade-off. We choose 0.5 to informally balance the two. Constants close to zero will yield high bias because $r_{n}=0$ corresponds to difference in means. Constants close to one result in high variance. To see this, suppose each cluster is a ball with homogeneous radius $R$. Setting the constant to 1 implies $r_{n}=R$, so the only unit in any cluster that is well surrounded with probability one is the centroid. A unit located slightly north of the centroid finds its $r_{n}$-neighborhood intersecting at least two clusters, so the chance that it is well surrounded, and hence not excluded, is substantially lower. In contrast, suppose $r_{n}=0.5R$, and define a cluster $C_{j}$’s “interior” as the ball centered at $m_{j}$ with radius $R/2$. Then all units in the interior are well surrounded with probability one because their $r_{n}$-neighborhoods are contained within $C_{j}$.
@@
$\Lambda_{i}=\big{\{}\ell\in\mathcal{N}_{n}\colon\max_{j\in\{1,\ldots,k_{n}\}}\,\lvert C_{j}\cap\mathcal{N}(i,r_{n})\rvert\,\lvert C_{j}\cap\mathcal{N}(\ell,r_{n})\rvert>0\big{\}},$ (3)
the set of units $\ell$ for which some cluster intersects the $r_{n}$-neighborhood of $\ell$ and $i$. These can be thought of as the units “most potentially correlated” with $i$. Define $A_{ij}(1)=\bm{1}\{j\in\Lambda_{i}\}$, $A_{ij}(2)=\bm{1}\{j\in C_{c(i)}\}$, and $\hat{Z}_{i}=(T_{1i}(Y_{i}-\hat{\mu}_{1}))/p_{1i}-(T_{0i}(Y_{i}-\hat{\mu}_{0}))/p_{0i}$. The variance estimator is
@@
Notice that $\hat{\sigma}^{2}(2)$ is the conventional cluster-robust variance estimator (e.g. Baird et al., 2018), which only accounts for within-cluster dependence. The estimator $\hat{\sigma}^{2}(1)$ is analogous to that of Leung (2022b) and additionally accounts for cross-cluster dependence since $C_{c(i)}\subseteq\Lambda_{i}$. In § 4, we discuss the advantages of taking the larger of the two.
@@ 4 Asymptotic Theory
We next derive bounds on the rates of convergence of our estimator and difference in means. We then characterize the asymptotic distribution of our estimator and prove that the variance estimator is asymptotically conservative.
@@
For all $t\in\{0,1\}$, if $d_{t}=1$ ($d_{t}=0$), then $p_{t}>0$ ($p_{t}<1$).
As discussed in Remark 1, this ensures overlap or positivity. For the next two theorems, abbreviate the estimand $\theta^{*}(d_{1},d_{0},p_{1},p_{0})$ as $\theta^{*}$. Recall the asymptotic order notation from the end of § 1.
@@
The result establishes a bias-variance trade-off in $k_{n}$. The $k_{n}^{-1/2}$ term is the contribution of (the square root of) the variance since the effective sample size in a CRT is the number of clusters $k_{n}$. The asymptotic bias is order $r_{n}^{-\gamma}$. Intuitively, if $r_{n}$ is large, then the expected outcome of a unit $i$ assigned treatment $d_{t}$ and well surrounded by clusters assigned to saturation level $p_{t}$ should well approximate ${\bf E}^{*}_{p_{t}}[Y_{i}(d_{t},\bm{D}_{-i})]$, corresponding to lower bias. The bias decreases at a faster rate for larger $\gamma$ since this corresponds to a lower degree of spatial interference.
The requirement $k_{n}\precsim n/\xi_{n}$ is mild, and typically we would have $k_{n}\prec n/\xi_{n}$. In the increasing-domain case where $\xi_{n}=1$, necessarily $k_{n}\precsim n/\xi_{n}$ since $k_{n}\leq n$, and usually the number of clusters is of smaller order than the population size. In the infill-increasing case, if $k_{n}$ grows faster than $n/\xi_{n}$, which is the volume of the study region under Assumption 2, then we would be increasingly subdividing the region into smaller clusters of shrinking volume, analogous to having more clusters than units.
Denote the difference-in-means estimator of $\theta^{*}$ by $\hat{\theta}^{+}$, which corresponds to setting $r_{n}=0$ in the definition of $\hat{\theta}$. Let $\hat{\theta}_{D}^{+}$, $\hat{\theta}_{T}^{+}$, and $\hat{\theta}_{O}^{+}$ be the difference-in-means estimates of $\theta_{D}^{*}$, $\theta_{T}^{*}$, and $\theta_{O}^{*}$ defined in § 2.3.
@@
(a) Under Assumptions 16, $\lvert\hat{\theta}^{+}-\theta^{*}\rvert\precsim(k_{n}\xi_{n}/n)^{1/d}+k_{n}^{-1/2}$. (b) Suppose the design satisfies Assumption 4 for $p_{1}\in(0,1)$, $p_{0}=0$. There exist a metric space and sequence of units, clusters, and potential outcomes satisfying Assumptions 13 and 5 such that $\lvert\hat{\theta}_{Q}^{+}-\theta_{Q}^{*}\rvert\succsim(k_{n}\xi_{n}/n)^{1/d}+k_{n}^{-1/2}$ for any $Q\in\{D,T,O\}$.
Part (a) provides an upper bound on the rate. Part (b) shows that the rate is tight for several illustrative cases. Recall that setting $p_{0}=0$ corresponds to the “pure control” baseline used in the estimands of Hayes and Moulton (2017).
The $k_{n}^{-1/2}$ term in the bound is (the square root of) the variance contribution, as in Theorem 2, while $(k_{n}\xi_{n}/n)^{1/d}$ is the bias contribution. The latter is notably worse than that of $\hat{\theta}$ because a reduction in the degree of interference $1/\gamma$ has no effect on the rate. The reason is that units situated at cluster boundaries may be directly proximate to clusters assigned to different treatment arms, and as shown in the proof, the share of such units can be of order $(k_{n}\xi_{n}/n)^{1/d}$.
These results demonstrate that excluding units in the manner of $\hat{\theta}$ can significantly reduce the asymptotic order of the bias relative to difference in means. While excluding units may come at the cost of efficiency, there is no increase to the asymptotic order of the variance because the variances of both estimators scale not with the number of units but with the number of clusters. Hence, the efficiency loss is second-order relative to the potential reduction in bias.
@@
Note that $\sigma_{n}^{2}\precsim 1$ by the proof of Theorem 2.
@@
The first result (5) centers the estimator at $\bar{\theta}$, the probability limit of $\hat{\theta}$. This is not the estimand of interest since there is an asymptotic bias $\lvert\bar{\theta}-\theta^{*}\rvert$ by Theorem 2. To use the normal limit to justify the validity of conventional CIs for $\theta^{*}$, the second result (6) requires “undersmoothed designs” in which the number of clusters is of smaller order than the optimal rate discussed in the next section. This ensures that the asymptotic bias is small. It is analogous to nonparametric regression where rate-optimal tuning parameter choices result in asymptotic bias, so conventional CIs require undersmoothing. We discuss practical choices of $k_{n}$ in § 5. In § SA.2.3, we discuss “bias-aware” inference, an alternative to undersmoothing.
@@
Suppose $1\prec k_{n}\prec n/\xi_{n}$ and $\sigma_{n}^{2}\succsim 1$. Under Assumptions 16, $\hat{\sigma}^{2}(t)=\sigma_{n}^{2}+\mathcal{B}_{n}+o_{p}(1)$ for $t\in\{1,2\}$ and some sequence of non-negative constants $\{\mathcal{B}_{n}\}_{n\in\mathbb{N}}$. Hence, both $\hat{\sigma}^{2}(1)$ and $\hat{\sigma}^{2}(2)$ are asymptotically conservative.
The cluster-robust variance estimator $\hat{\sigma}^{2}(2)$ has the advantage of being non-negative in finite sample, unlike $\hat{\sigma}^{2}(1)$ which is a truncation estimator (Andrews, 1991, p. 823). On the other hand, $\hat{\sigma}^{2}(2)$ only accounts for within-cluster dependence, whereas $\hat{\sigma}^{2}(1)$ also accounts for cross-cluster dependence in the definition $A_{ij}(1)$. As shown in the proof, this dependence vanishes, so $\hat{\sigma}^{2}(1)$ is a valid estimator. However, the dependence vanishes at the slow rate $(k_{n}\xi_{n}/n)^{1/d}$, the same order as the bias of difference in means. Thus in smaller samples, ignoring second-order terms may result in anti-conservativeness. By taking the larger of the two estimators, we obtain the benefits of both. See § SA.3 for a comparison of $\hat{\sigma}^{2}$ with other variance estimators in the literature.